OpenStudy (anonymous):

Double checking. 1.The product of $25 and "n" of club memberships is greater than or equal to $500. $25(n)>=$500 2.The distance to school, "s" is 1/5 of a mile more than the distance of "c" to the Community Center swimming pool. s=1/5+c 3.The simple interest earned on a principal of $300 at an annual interest rate of "x" percent is less than or equal to $72. Ok, I don't know how to do that one....help me?

OpenStudy (akshay_budhkar):

first one is correct

OpenStudy (akshay_budhkar):

second one is also correct :)

OpenStudy (anonymous):

Akshay, is it safe to accept this?: has sent you an audio/video chat request. Click here to accept it or simply ignore this message.

OpenStudy (anonymous):

From kushawa somethingfish.

OpenStudy (akshay_budhkar):

well it doesnt work anyways!! LOL i still have to configure it Better ignore it

OpenStudy (anonymous):

Kusha said accept.....idk...do I say you didn't figure it out?

OpenStudy (akshay_budhkar):

say you accepted it but it is not working and if you do not want to chat to him all together better ignore him

OpenStudy (anonymous):

Ok, I don't know him. So, I'll go with your 1st option. :)

OpenStudy (akshay_budhkar):

ok the last one now

OpenStudy (akshay_budhkar):

simple interest is (P*n*r)/100 you know this?

OpenStudy (anonymous):

No, I did not. Can you explain the parenthesis, *'s and so on as to how you got it?

OpenStudy (akshay_budhkar):

Well its the formula Simple interest= Principle*number of years*rate of interest/100 i do not know how we get it but just memorized it

OpenStudy (aravindg):

akshay u there???

OpenStudy (anonymous):

Oh...ok. Thank you Akshay. :D Could you wait a bit while I post another one on the left? .........unless it's inconveniencing you then you don't have to...-______-

OpenStudy (anonymous):

Oh and thank you for your advice regarding, Kusha. I logged out and I'm going to be here now. :)

OpenStudy (akshay_budhkar):

LOL

OpenStudy (aravindg):

akshay can u help me its urgent???

OpenStudy (akshay_budhkar):

i am actually very busy! we are having a festival tomorrow! and i am dead busy

OpenStudy (aravindg):

5 min plzzzzzzz

OpenStudy (anonymous):

Oh....ok. I apologize, Akshay. Thank you for your time and help though!!! xD You're awesome!!!! Good luck with your festival, Akshay. I wish you the best!!!!!! :D

OpenStudy (akshay_budhkar):

yea arvind ask me

OpenStudy (akshay_budhkar):

nothing to apologize jessica.. ask me here i will answer it

OpenStudy (anonymous):

?! Um..ok. :D I'm not sure how to go about writing this: The square of the length "c" of the hypotenuse of a right triangle is equal to four squared plus three squared. (Write an equation or an inequality to model the real-life situation)

OpenStudy (akshay_budhkar):

c^2=4^2+3^2

OpenStudy (anonymous):

Can you explain how you got all that? The same way you explained the 3rd question?

OpenStudy (akshay_budhkar):

see square of c is c^2

OpenStudy (akshay_budhkar):

is equal to will give you equal to sign

OpenStudy (akshay_budhkar):

four squared= 4^2

OpenStudy (akshay_budhkar):

plus is +

OpenStudy (akshay_budhkar):

and three squared is 3^2

OpenStudy (aravindg):

tanx +tan(x+pi/3) +tan(x+2pi/3)=3

OpenStudy (anonymous):

Oh....ok. That makes more sense now. :) Thank you, Akshay. :D

OpenStudy (akshay_budhkar):

oh triangle trignometry!! the worst part of maths

OpenStudy (aravindg):

then a)tan 2x=1 b)tan 3x=1 c)tan^2 x=1 d)tan^3x=1

OpenStudy (aravindg):

helppp

OpenStudy (akshay_budhkar):

umm.. i am sorry man!! i dont know Why not open tan(x+pi/3) using formula?